OpenStudy (goformit100):

The sides of a triangle are three consecutive integers and its inradius is four units. Determine the circumradius.

OpenStudy (goformit100):

@Mertsj Help me sir

OpenStudy (anonymous):

inradius: \[r = \frac{1}{2}\sqrt{\frac{(b + c - a)(a+c-b)(a+b-c)}{a + b + c}}\] given r = 4, let a = n, b = n+1, c = n+2: \[8=\sqrt{\frac{(n+3)(n+1)(n-1)}{3(n+1)}}\] \[8=\sqrt{\frac{n^2+2n-3}{3}}\] \[\sqrt{64}=\frac{\sqrt{n^2+2n-3}}{\sqrt{3}}\] \[\sqrt{192}=\sqrt{n^2+2n-3}\] Find n, then use: \[r=\frac{abc}{4sR}\] where s is the semiperimeter \(\frac{1}{2}(a+b+c)\) and R is the circumradius

OpenStudy (goformit100):

Thank you Madam